Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 16     Entry time: 08 Feb 2024, 23:14
Author: Tokushi Sato 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 1-350 at 12.45 keV, 1700 uJ. Stable delivery

Optical laser delivery

  • None

Achievements/Observations

  • fixing many issues with the cube and helium leaks
  • collected many quality diffraction data for both jet with and without samples
  • data on Nendu mutant (~2.5 hours of data collection) with two different conditions

Issues

  • AGIPD data aggregator stopped responding and it took about 1 h to be fixed by DOC, which interrupted the data collection

  • some strange streak diffraction appears on JF4M and this has to be removed by scanning X- and Y- nozzle motors carefully
ELOG V3.1.4-7c3fd00